A particle with a charge of -1.24x10^-8 C is moving with instantaneous velocity (1.49x10^4m/s) î +(-3.85x10^4m/s) ĵ.​ What is the force exerted on this particle by a magnetic field.

(a) (1.40T)î+ (-2.30T)

(b) (-1.25T)î+ (2.00T)ĵ + (-3.50T)

1
Expert's answer
2021-06-24T16:13:33-0400

(a) "\vec F=q[\vec v, \vec B]=-1.24\cdot10^{-8}\cdot[1.49\cdot10^4\vec i-3.85\cdot10^4\vec j, 1.4\vec i-2.3\vec k]="


"=-0.0011\vec i-0.00042\vec j-0.00067\vec k" ; "|\vec F|=0.0014\ (N)"


(b) "\vec F=q[\vec v, \vec B]=-1.24\cdot10^{-8}\cdot[1.49\cdot10^4\vec i-3.85\cdot10^4\vec j, -1.25\vec i+2\vec j-3.5\vec k]="


"=-0.0017\vec i-0.00065\vec j+0.00023\vec k" ; "|\vec F|=0.0018\ (N)"